Two friends, Al and Jo, have a combined mass of 168 kg. At an ice skating rink they stand close together on skates, at rest and facing each other, with a compressed spring between them. The spring is kept from pushing them apart because they are holding each other. When they release their arms, Al moves off in one direction at a speed of 0.75 m/s, while Jo moves off in the opposite direction at a speed of 1.5 m/s. Assuming that friction is negligible, find Al's mass. kg
Answer #1

here,

let the mass of Al and JO be m1 and m2

m1 + m2 = 168 kg ....(1)

and

using conservation of momentum

m1 * v1 + m2 * v2 = 0

m1 * 0.75 = m2 * 1.5

m1 = 2 * m2 ...(2)

from (1) and (2)

m2 = 56 kg

m1 = 112 kg

the Al's mass is 112 kg